A man of mass 60 kg jumps from a trolley of mass 20 kg standing on smooth surface with absolute velocity 3m/s. Find velocity of trolley and total energy produced by man.

Grade:11

1 Answers

Swapnil Saxena
12 years ago

Since the trolley and man were initially at rest Their initial momenta is 0.

Once the man jumps and attains a velocity of 3m/s (with respect to ground) its momnetum is 60*3= 180kgm/s

Accor. to the law of conservation of momentum ,

0= 180+20v

v= 9m/s in the opposite direction.

Assuming that no deformation occur in the trolley ( deformation causes loss of energy) ,total energy that is produced by the man = 1/2(20*92+60*32 ) = 1/2((20*81)+(60*9)) = 1/2((1620)+(540)) = 1/2(2160)= 1080 J